TechMaster is an information technology (IT) consulting company offering services to small firms. TechMaster bills clients fob. Prepare an income statement for the company for August. TechMaster Income Statement For the Month Ended August 31
Answer #1
Wages Payable $85*885 =$75225
Wages Payable
opening Balance
Ending bal. 75225 Work in progress 75225
Work­ In­ Process= $47*875
Work­ In­ Process
Wages payable 75225 Cost of services billed 116350
Applied Service O.H 41125
Cost of services billed =41125-37000 = 4125
Cost of services billed
Work in progress 116350 Service OH control 4125
Closing Balance 112225
Service Overhead control
Cost of services billed 37000 Closing balance 37000
   Applied Service Overhead
Cost of service billed 41125 Work in progress 41125
TechMaster
Income statement
For the month ending August 31
Particulars Amount
Sales ($220*885) $194,700
Cost of service billed 116350
Less:Over applied service Overhead -4125 $112,225
Gross margin $82,475
marketing and administrative costs $57,000
Operating Profit $25,475